A living room has a circular rug. According to a small tag, the rug has an area of 3.14 square meters. What is the rug's circumference?
Use 3.14 for π.
meters

Answer by rfer(16322)

A=3.14r^2
3.14=3.14r^2
3.14/3.14=r^2
r^2=1
r=1
-----------------
3.14*2=6.28 meters